# Align Your Intents: Offline Imitation Learning via Optimal Transport

### v1

@@ Abstract
Offline reinforcement learning (RL) addresses the problem of sequential decision-making by learning optimal policy through pre-collected data, without interacting with the environment. As yet, it has remained somewhat impractical, because one rarely knows the reward explicitly and it is hard to distill it retrospectively. Here, we show that an imitating agent can still learn the desired behavior merely from observing the expert, despite the absence of explicit rewards or action labels. In our method, AILOT (Aligned Imitation Learning via Optimal Transport), we involve special representation of states in a form of intents that incorporate pairwise spatial distances within the data. Given such representations, we define intrinsic reward function via optimal transport distance between the expert’s and the agent’s trajectories. We report that AILOT outperforms state-of-the art offline imitation learning algorithms on D4RL benchmarks and improves the performance of other offline RL algorithms in the sparse-reward tasks.
Figure 1: AILOT: Aligned Imitation Learning via Optimal Transport. Principal diagram for intents alignment. Left: Stage I: projection into intent space (denoted by $\psi$ in text) (only first $3$ principal components are shown). Middle: Stage II: computation of intrinsic rewards for offline RL, where $s_{i}^{a}$ and $s_{j}^{e}$ are the expert’s and the agent’s states, $P$ is the optimal transport transitions matrix; and the corresponding reward $r(s_{i}^{a})$ is a scaling transform of the product $\sum_{j}P_{ij}C_{ij}$ with some cost function $C$ defined on the intents pairs. Right: squared norm vs. steps count in the same trajectory for the states and the intents differences; It demonstrates that distance between intents is proportional to the total path length (steps count) between the states (AntMaze example is shown).
@@ 1 Introduction
Over the past years, offline learning has remained both the most logical and the most ambitious avenue for the development of RL. On the one hand, there is an ever-growing reservoir of sequential data, such as video, becoming available for training the decision-making RL agents. On the other hand, these immense data remain largely unlabeled and unstructured for gaining any valuable guidance in the form of rewards or action labels, stimulating the development of unsupervised and self-supervised methods Singh et al. (2020); Li et al. (2023); Sinha et al. (2022); Yu et al. (2022)
Drawing inspiration from the triumphant strategies employed in language-based foundational models, leveraging the vast reservoir of web-based data for offline RL indeed emerges as a promising avenue. However, the unresolved challenge persists in efficiently incorporating offline data into RL frameworks and processes Levine et al. (2020). Offline RL faces several challenges, such as distributional shift, slow convergence when the labels are unavailable, and the absence of known rewards which should be purposely designed for each separate task independently Kumar et al. (2021); Fujimoto et al. (2019).
@@
A potential remedy to these challenges is found in Imitation Learning (IL), where the explicit reward functions are not needed. Instead, an imitating agent is trained to replicate the behavior of the expert. Behavior cloning (BC) Ross and Bagnell (2010) frames the IL problem akin to classical supervised learning, seeking to maximize the likelihood of the actions provided under the learner’s policy. Despite working well in simple environments, BC is prone to accumulating errors in states, coming from different distributions other than experts. Distribution matching is another promising IL paradigm, where approaches such as distribution correction estimation (”DICE”), including SMODICE Ma et al. (2022b), attempt to match state-occupancy measures between the imitator’s and the expert’s policies. Another notable approach is DEMODICE Kim et al. (2022), which uses state-action occupancy measures. One significant limitation of these methods is their requirement for a non-zero overlap between the supports of the agent’s performance dataset and the expert data.
@@
Another promising avenue involves Computational Optimal Transport, a domain that has garnered significant popularity for addressing diverse machine learning tasks Peyré et al. (2019). Its applications span from domain adaptation to generative modeling Salimans et al. (2018); Rout et al. (2021); Korotin et al. (2022). It has proven to be efficient in weakly-supervised tasks Bespalov et al. (2022) and even for learning with one target sample and unlabelled source data Bespalov et al. (2020). To alleviate the necessity for manual reward engineering, Luo et al. (2023) leverages optimal transport theory. The approach automates the assignment of reward labels by establishing an optimal coupling between a select few high-quality expert demonstrations and the trajectories of the learning agent.
The absence of reward labels is not the sole challenge; obtaining access to expert actions can also be problematic. One potential remedy is to create a simulator Krylov et al. (2020); Saboo et al. (2021) that enables access to both actions and rewards. However, adopting such an approach entails substantial additional work. Consider a scenario where a thousand demonstrations of expert trajectories are available from videos, yet no rewards or action labels accompany them. In this work, we eliminate the requirement for either knowing the rewards or the action labels of the expert. By utilizing the Intention-Conditioned Value Function (ICVF) Ghosh et al. (2023), we showcase the potential to redefine the offline Imitation Learning (IL) problem as expert-guided. This guidance is achieved by contracting intent representations through the optimal transport. This framework allows for the complete discarding of the action labels, with the imitation now being centered around the proximity to the expert intentions. Given that an intention is a task-agnostic property, the learned representations prove to be sufficiently discriminative, providing valuable insights into the environment without relying on specific action or reward labels.
Our contribution is as follows:
  • We introduce a new intrinsic reward function that depicts the dynamics in the environment and an offline imitation learning algorithm based on it.
@@
  • We report extensive comparison with previous works. Despite not having access to the expert’s action labels and ground truth rewards, we outperform the state-of-the-art models in the majority of benchmark datasets and show that our approach enables custom imitation even when agent’s data is a mix of random policies.
@@ 2 Related Works
In this study, we broaden the application of offline Reinforcement Learning (RL) to datasets that lack both the rewards and the action labels. Despite all the research effort on learning the intrinsic rewards for RL, most works assume either online RL setting Brown and Niekum (2019); Yu et al. (2020); Ibarz et al. (2018) or the unrealistic setup of possessing some annotated prior data. Moreover, the problem of learning and exploration in sparse-reward environments can be deemed as solved for online RL Li et al. (2023); Eysenbach et al. (2018); Lee et al. (2019); whereas, only a few works exist that consider the offline goal-conditioned setting with reward-free data Park et al. (2023); Zheng et al. (2023); Wang et al. (2023). The goal of our work is to extract guidance from observing the expert by finding optimal alignment of information-rich representations of imaginary goals between the expert and the agent in a shared latent space.
We build upon the approach introduced in Luo et al. (2023), extending it by finding a representative latent space. In Luo et al. (2023), the authors address the entropy-regularized Optimal Transport (OT) problem using Sinkhorn’s algorithm, focusing on expert and agent trajectories. By estimating the optimal coupling between these trajectories, every trajectory in the offline dataset can be annotated by intrinsic rewards.
A clear advantage of our method is its independence from the action and the reward labels, distinguishing it from the prior works Ho and Ermon (2016); Kim et al. (2022); Garg et al. (2021); Reddy et al. (2019); Pomerleau (1991); Yu et al. (2022). Additionally, the method seamlessly integrates with various downstream RL algorithms, providing flexibility in selecting the most suitable training approach. It is also essential to acknowledge a limitation of Luo et al. (2023): the method does not guarantee that geometry of unlabeled states distribution will be discriminative enough to provide right non-noisy reward signal through optimal transport. An alternative research direction, Calibrated Latent gUidancE (CLUE) Liu et al. (2023), introduces a parallel approach for deriving intrinsic rewards. In this study, the authors employ a conditional variational auto-encoder trained on both expert and agent transitions and compute the distance between the collapsed expert embedding and the agent trajectory. The expert embedding might not collapse into a single point in a multimodal expert dataset, requiring clustering to handle different skills. Unlike the CLUE method, our approach doesn’t need state-action labeled data, eliminating the need for action annotations.
@@ 3.0.1 Problem Formulation
A standard Reinforcement Learning problem is defined as a Markov Decision Process (MDP) with tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},p,r,\rho_{0},\gamma)$, where $\mathcal{S}\subset\mathbb{R}^{n}$ is the state space, $\mathcal{A}\subset\mathbb{R}^{m}$ is the action space, $p$ is a function describing transition dynamics in the environment $p$: $\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{P}(\mathcal{S})$, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is a predefined extrinsic reward function, $\rho_{0}$ is an initial state distribution and $\gamma\in(0,1]$ is the discount factor. The objective is to learn policy $\pi_{\theta}(a|s):\mathcal{S}\rightarrow\mathcal{P}(\mathcal{A})$ that maximizes discounted cumulative return $\mathbb{E}[\sum^{\infty}_{t=0}\gamma^{t}r(s_{t},a_{t})]$. In contrast, offline RL assumes access to a pre-collected static dataset of transitions $\mathcal{D}=\{(s_{i},a_{i},s^{\prime}_{i},r_{i})\}^{n}_{i=1}$ while strictly prohibiting interaction with the environment. In this study, we assume access to a dataset of transitions without reward labels $\mathcal{D}_{a}=\{(s_{i},a_{i},s^{\prime}_{i})\}^{n}_{i=1}$, collected from $\mathcal{M}$, and a limited number of ground truth demonstrations from the expert policy without reward and action labels $\mathcal{D}_{e}=\{(s_{i},s^{\prime}_{i})\}^{m}_{i=1}$. These demonstrations are also collected from the same MDP ($\mathcal{M}$) and we assume that trajectories from $\mathcal{D}_{e}$ have high cumulative return. The primary objective is to determine a policy that closely emulates the behavior of the expert and thereby maximises the cumulative return.
3.0.2 ICVF Pretraining
Using pre-collected large unlabeled datasets in the form of prior data empowers agents to generalize their learned behaviors. Even in the absence of ground truth actions or rewards, agents can still acquire valuable insights into the environment dynamics and learn useful features. Recent work in the generalization of the value function, known as the Intention-Conditioned Value Function (ICVF), introduces the concepts of intentions and outcomes to replace traditional notions of actions and rewards Ghosh et al. (2023). ICVF fundamentally constructs a successor representation Dayan (1993), allowing the recovery of spatial correspondences within data based solely on observations. ICVF is task-agnostic, and the learned embeddings demonstrate sufficient discriminative capabilities to guide learning agents through intrinsic rewards toward expected behavior. Formally, considering two states from an offline dataset $s,s_{+}\in\mathcal{D}=\mathcal{D}_{a}\bigcup\mathcal{D}_{e}$, ICVF is defined as:
$V(s,s_{+},z)=\sum_{t\geq 0}\gamma^{t}\mathbb{E}_{P_{z}}\left[\mathbb{I}[s_{t}=s_{+}]-1\,\big{|}\,s_{0}=s,z\right].$ (1)
Let $P_{\mathcal{D}}(s_{t+1}|s_{t})$ denote the empirical transition probability in the dataset $\mathcal{D}$. The corresponding transition matrix of the Markov process is
$P_{z}(s_{t+1}|s_{t})=\frac{\pi(z|s_{t+1},s_{t})}{\pi(z|s_{t})}P_{\mathcal{D}}(s_{t+1}|s_{t}).$ (2)
ICVF estimates an average path length between states $s$ and $s_{+}$ (with a negative sign), conditioning on guidance of a latent variable $z$, defined as an intent of another state $\psi(s_{z})$. The intents policy $\pi(z|s_{z},s)$ implicitly (by means of implicit Q-learning method Kostrikov et al. (2021)) minimizes total step count between $s$ and $s_{z}$ from the same trajectory:
$\min_{\pi,\psi}\mathbb{E}_{\mathcal{D}}(-1)\log\pi(\psi(s_{z})|s_{z},s)V(s,s_{z},\psi(s_{z})).$ (3)
As a part of ICVF procedure, the value function is parametrized by three neural networks in the following decomposed format:
@@ 3.0.2 ICVF Pretraining
We observed that the $\psi$ mapping itself (without $\phi$ and $T$) serves as a good estimate for the distances between the states of the environment. In the Experiments section below, we will empirically support this, showing that squared distance $d^{2}(s_{t+k},s_{t})=\|\psi(s_{t+k})-\psi(s_{t})\|^{2}$ is roughly a linear function of the steps count between the states ($k$).

### v2

@@ Abstract
Offline Reinforcement Learning (RL) addresses the problem of sequential decision-making by learning optimal policy through pre-collected data, without interacting with the environment. As yet, it has remained somewhat impractical, because one rarely knows the reward explicitly and it is hard to distill it retrospectively. Here, we show that an imitating agent can still learn the desired behavior merely from observing the expert, despite the absence of explicit rewards or action labels. In our method, AILOT (Aligned Imitation Learning via Optimal Transport), we involve special representation of states in a form of intents that incorporate pairwise spatial distances within the data. Given such representations, we define intrinsic reward function via optimal transport distance between the expert’s and the agent’s trajectories. We report that AILOT outperforms state-of-the art offline imitation learning algorithms on D4RL benchmarks and improves the performance of other offline RL algorithms by dense reward relabelling in the sparse-reward tasks.
@@ 1 Introduction
Over the past years, offline learning has remained both the most logical and the most ambitious avenue for the development of RL. On the one hand, there is an ever-growing reservoir of sequential data, such as videos, becoming available for training the decision-making of RL agents. On the other hand, these immense data remain largely unlabeled and unstructured for gaining any valuable guidance in the form of a tractable objective for learning the underlying policy, stimulating the development of unsupervised and self-supervised methods (Singh et al., 2020; Li et al., 2023; Sinha et al., 2022; Yu et al., 2022; Eysenbach et al., 2018; Park et al., 2023b).
Following the success of language-based foundational models, the offline RL community have also begun to leverage the weakly-labeled data. However, incorporating such offline data into RL frameworks efficiently remains a challenge (Levine et al., 2020)). Several factors, such as distribution shift, slow convergence when the labels are missing, and the absence of known rewards or task-specific objectives, hinder the development of offline RL (Kumar et al., 2021; Fujimoto et al., 2019).
@@
A potential remedy to these issues is found in Imitation Learning (IL), where the explicit reward function is not needed. Instead, an imitating agent is trained to replicate the behavior of the expert. Behavior cloning (BC) (Ross & Bagnell, 2010) frames the IL problem akin to classical supervised learning, seeking to maximize the likelihood of the actions provided under the learner’s policy. Despite working well in simple environments, BC is prone to accumulating errors in states, coming from different distributions other than that of the expert. Other notable methods include Inverse RL (Torabi et al., 2018; Zolna et al., 2020) and action pseudo-labeling (Kumar et al., 2020), however they are rarely used in practice due to the introduced overhead. Distribution matching is another promising IL paradigm, where the approaches such as DIstribution Correction Estimation ("DICE"), including SMODICE (Ma et al., 2022b), attempt to match the state-occupancy measures between the imitator’s and the expert’s policies. However, these methods posses several limitations: 1) they require a non-zero overlap between the supports of the agent and the expert and 2) they mostly use KL-divergence (or, generally, f-divergences), which ignores the underlying geometry of the spaces where the distributions are defined.
@@
Another line of development is associated with Computational Optimal Transport, a domain that has garnered significant popularity for addressing diverse machine learning tasks (Peyré et al., 2019). Its applications span from domain adaptation to generative modeling (Salimans et al., 2018; Rout et al., 2021; Korotin et al., 2022), including weakly-supervised and unsupervised methods (Bespalov et al., 2020; 2022). Optimal transport theory was proposed to alleviate the necessity for manual reward engineering (Luo et al., 2023) via establishing an optimal coupling between a few high-quality expert demonstrations and the trajectories of the learning agent, which proved to be useful for Imitation Learning tasks (Haldar et al., 2022; 2023). However, performing OT matching in the high-dimensional unstructured space is difficult, thus limiting the approach to low-dimensional tasks.
The absence of reward labels is not the sole challenge; obtaining access to the expert actions can also be problematic. When no rewards or action labels accompany the abundant demonstrations, one potential remedy is to create a simulator (Krylov et al., 2020; Saboo et al., 2021) that enables access to both actions and rewards. However, adopting such an approach entails substantial additional work. Here, we eliminate the requirement for knowing the rewards or the action labels of the expert altogether. We propose to map the initial state space to the space of intentions, where there is some high-level semantic imaginary goal that the agent needs to achieve to imitate the expert. Aligning intents of the agent with those of the expert via Optimal Transport represents a new pathway for training efficient offline RL models.
Our contribution are as follows:
  • A new intrinsic dense reward relabelling algorithm that ‘comprehends’ the demonstrated dynamics in the environment, on top of which any offline RL method can be used.
  • We outperform the state-of-the-art models in the majority of Offline RL benchmarks. We do so even without knowing the expert’s action labels and the ground truth rewards.
@@
  • We report extensive comparison with previous works. We show that our approach enables custom imitation even if the agent’s data are a mix of random policies.
@@ 2 Related Works
In this study, we broaden the application of offline Reinforcement Learning (RL) to datasets that lack both the rewards and the action labels. Despite all the research effort on learning the intrinsic rewards for RL, most works assume either online RL setting (Brown & Niekum, 2019; Yu et al., 2020; Ibarz et al., 2018) or the unrealistic setup of possessing some annotated prior data. Moreover, the problem of learning and exploration in sparse-reward environments can be deemed as solved for online RL (Li et al., 2023; Eysenbach et al., 2018; Lee et al., 2019); whereas, only a few works exist that consider the offline goal-conditioned setting with reward-free data (Park et al., 2023a; Zheng et al., 2023; Wang et al., 2023a). The goal of our work is to extract guidance from the expert by forcing optimal alignment of information-rich representations of imaginary goals between the expert and the agent in a shared distance-aware latent space.
We build upon the approach introduced in OTR (Luo et al., 2023), extending it by finding a representative distance-preserving isometry to the shared latent space and forcing alignment via optimal transport in this metric-aware space. In OTR, the authors estimate the optimal coupling between expert and agent trajectories, which enables rewards relabeling for agent’s state transitions by computed Wasserstein distance. However, performing optimal transportation matching in this way is sensitive to underlying cost function (e.g Cosine or Euclidean) and final results can vary drastically. Choosing right distance function requires additional knowledge of the environment and extensive search, limiting application to simple tasks.
An advantage of our method is in its ability to perform distribution alignment in the space, which captures temporally structured dependencies in the provided datasets, which is completely ignored in previous works. For example, a lot of methods focus on KL divergence, which is agnostic to the distance metric Such temporal structure induces spatial closeness between temporally similiar states.Additionally, the method seamlessly integrates with various downstream RL algorithms, providing flexibility in selecting the most suitable training approach.
Another recent method called Calibrated Latent gUidancE (CLUE) (Liu et al., 2023), introduces a parallel approach for deriving intrinsic rewards. In this study, the authors employ a conditional variational auto-encoder trained on both expert and agent transitions and compute the euclidean distance between the collapsed expert embedding and the agent trajectory. The expert embedding might not collapse into a single point in a multimodal expert dataset, requiring clustering to handle different skills. Unlike the CLUE method, our approach doesn’t need state-action paired labeled data, eliminating the need for action annotations.
@@ 3.1 Problem Formulation
A standard Reinforcement Learning problem is defined as a Markov Decision Process (MDP) with tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},p,r,\rho_{0},\gamma)$, where $\mathcal{S}\subset\mathbb{R}^{n}$ is the state space, $\mathcal{A}\subset\mathbb{R}^{m}$ is the action space, $p$ is a function describing transition dynamics in the environment $p$: $\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{P}(\mathcal{S})$, $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is a predefined extrinsic reward function, $\rho_{0}$ is an initial state distribution and $\gamma\in(0,1]$ is the discount factor. The objective is to learn policy $\pi_{\theta}(a|s):\mathcal{S}\rightarrow\mathcal{P}(\mathcal{A})$ that maximizes discounted cumulative return $\mathbb{E}[\sum^{\infty}_{t=0}\gamma^{t}r(s_{t},a_{t})]$. In contrast, offline RL assumes access to a pre-collected static dataset of transitions $\mathcal{D}=\{(s_{i},a_{i},s^{\prime}_{i},r_{i})\}^{n}_{i=1}$ while prohibiting additional interaction with the environment. In this study, we assume access to a dataset of transitions without reward labels $\mathcal{D}_{a}=\{(s_{i},a_{i},s^{\prime}_{i})\}^{n}_{i=1}$, collected from $\mathcal{M}$, and a limited number of ground truth demonstrations from the expert policy without reward and action labels $\mathcal{D}_{e}=\{(s_{i},s^{\prime}_{i})\}^{m}_{i=1}$. These demonstrations are collected from the same MDP ($\mathcal{M}$) and we assume that several trajectories from $\mathcal{D}_{e}$ have high cumulative return. The primary objective is to determine a policy that closely emulates the behavior of the expert and thereby maximises the cumulative return.
3.2 Reward Relabelling through Optimal Transport
Optimal Transport provides a reasonable and efficient way of comparing two probability measures supported on high-dimensional measure spaces. Given a defined cost function $c(\cdot,\cdot)$ over $\mathcal{S}\times\mathcal{S}$, the Wasserstein distance between the agent trajectory $\tau^{a}=\{s^{a}_{1},s^{a}_{2},\cdots,s^{a}_{T}\}\subset\mathcal{D}_{a}$ and the expert trajectory $\tau^{e}=\{s^{e}_{1},s^{e}_{2},\cdots,s^{e}_{T}\}\subset\mathcal{D}_{e}$ is defined as:
$W(\tau^{a},\tau^{e})=\min_{P\in\mathbb{R}^{T\times T}}\sum^{T}_{i=1}\sum^{T}_{j=1}c(s^{a}_{i},s^{e}_{j})P_{ij},$ (1)
satisfying the marginal constraints, with $p^{a},p^{e}$ being the state occupancy of the agent data and the expert dataset respectively:
$\sum^{T}_{i=1}P_{ij}=p^{e}(s_{j});\quad\sum^{T}_{j=1}P_{ij}=p^{a}(s_{i}).$ (2)
Then, a proxy reward function, given the optimal transport plan $P^{*}(\tau_{a},\tau_{e})$ to transform the agent state $s^{a}_{i}$ into the closest state of the expert, is:
$r_{i}=-\sum^{T}_{j=1}c(s^{a}_{i},s^{e}_{j})P^{*}_{ij}(\tau_{a},\tau_{e}).$ (3)
In this work, instead of computing the optimal transport between the initial states from the datasets, we propose to extract intentions of the expert in the form of an isometry map to the latent space, which enables a more accurate modelling of the distribution of the similiar states (based on temporal distances between them).
Figure 1: AILOT: Aligned Imitation Learning via Optimal Transport. Principal diagram for intents alignment of states in two different trajectories (in blue and red). Left: Stage I: projection into intent space (denoted by $\psi$ in text) (first $3$ principal components are shown). Middle: Stage II: computation of intrinsic rewards for offline RL, where $s_{i}^{a}$ and $s_{j}^{e}$ are the expert’s and the agent’s states, $P$ is the optimal coupling matrix; corresponding intrinsic reward $r(s_{i}^{a})$ is a scaling transform of the product $\sum_{j}P_{ij}C_{ij}$ with some cost function $C$ defined on the intents pairs. Right: squared norm vs. steps count in the same trajectory for the states and the intents differences; It demonstrates that distance between intents is proportional to the total path length (steps count) between the states (AntMaze example is shown).
3.3 Pretraining of Distance Preserving Representation
Using pre-collected large unlabeled datasets in the form of prior data enables learning diversified behaviors in the form of skills or general representations. Even in the absence of the ground truth actions or rewards, the policies can still acquire valuable insights about the dynamics in the environment. Several works tried to define valuable objective for learning from the unstructured data, including offline GCRL (Eysenbach et al., 2022; Park et al., 2023a) and skill discovery (Eysenbach et al., 2018; Sharma et al., 2019). Finding general mapping $\psi:\mathcal{S}\rightarrow\mathbb{R}^{d}$ which is a lossless compression of a state was addressed in several formulations including successor features (SFs) (Dayan, 1993; Barreto et al., 2017; Borsa et al., 2018) and Forward-Backward (FB) representations (Touati & Ollivier, 2021) with their extensions to the generalized value functions (Hansen et al., 2019; Touati et al., 2022; Ghosh et al., 2023; Bhateja et al., 2023). In the same vein, we learn successor features but enforce additional constraint on temporal structure. Finding optimal temporal goal-conditioned value function $V^{*}$ (Wang et al., 2023b) which depicts the minimal number of steps, required to arrive at certain goal $g$ from state $s$ to $s_{+}$ can be viewed as goal conditioned RL maximizing $r(s,g)=-\mathbbm{1}(s\neq g)$. ICVF (Ghosh et al., 2023) exchanges $g$ with intent $z$, generalizing previous definition of optimal goal-conditioned value function as
$V^{*}(s,s_{+},z)=\mathbb{E}_{s_{t+1}\sim P_{z}(\cdot|s_{t})}\left[\sum_{t\geq 0}-\mathbbm{1}(s_{t}\neq s_{+})\bigg{|}s_{0}=s\right],$ (4)
where intent $z$ completely specifies the state-occupancy dynamics through the transition function $P_{z}$. In practice, $V$ is parametrized by three neural networks with $\phi,\psi:\mathcal{S}\to\mathbb{R}^{d}$ and $T:\mathbb{R}^{d}\to\mathbb{R}^{d\times d}$ being a transition matrix, incorporating all possible transition dynamics of each $z$:
@@ 3.3 Pretraining of Distance Preserving Representation
During training, the following temporal distance loss is minimized:
$\mathcal{L}^{2}_{\tau}(-\mathbbm{1}(s\neq s_{+})+\gamma\bar{V}(s^{\prime},s_{+},z)-V(s,s_{+},z)),$ (6)
with $\mathcal{L}^{2}_{\tau}(x)=|\tau-\mathbbm{1}(A<0)|x^{2}$, $A=r_{z}(s)+\gamma\bar{V}(s^{\prime},s_{+},z)-\bar{V}(s,s_{+},z)$ being an expectile loss (Kostrikov et al., 2021), $A$ is an advantage of acting according to $z$, and $\bar{V}$ is a ‘delayed’ copy (target network) of $V$. Such a representation estimates an average path length between the states $s$ and $s_{+}$, conditioning on guidance of all possible choices of $z$. The pretraining procedure enables a downstream extraction of learned representations $\psi,\phi$ useful for the other tasks. In our experiments, we set intent of a state $s$ as $z(s)=\psi(s)$ and $r_{z}(s)=-\mathbbm{1}(s\neq s_{z})$ with a randomly sampled future state $s_{z}$ from the same trajectory. The $\psi$ mapping itself (without $\phi$ and $T$) serves as a good estimate for the distances between the states of the environment. In the Experiments section below, we will empirically support this, showing that a squared distance $d(s_{t+k},s_{t})=\|\psi(s_{t+k})-\psi(s_{t})\|^{2}$ is roughly a linear function of the steps count between the states $k$, which maps temporally similar states in the initial state space to the spatially closest in the intent space. Thus, applying optimal transport tools in this space makes it robust to noise or inaccuracies in the initial state space of the environment. Extracting the intents from the pretrained general-purpose value function enables capturing actual behavior from the trajectory.
We also provide a theoretical confirmation of the convergence of extracted intents to the temporal distance function between states (ref. A.1), showing explicitly that if the intentions of states $s$ and $s_{+}$ converges to each other, then the function $V(s,s_{+})$ approaches zero in the limit.